the population standard deviation of the lifetime of washing machines is estimated to be 900 hours, how large a sample must be taken in order to be 90% confident that the margin of error will not exceed 100 hours?

Select one:

a. 219.18

b. 220

c. 14.8

d. 15
Answer #1

= 900. Given that The Standard deviation for 90% of confidence, 24/2 = 1.645. E = 24/2 h Margin of error = 1.645 (900) > 1480